Vernon Company makes a product that sells for $33 per unit. The company pays $24 per unit for the variable costs of the product and incurs annual fixed costs of $72,900. Vernon expects to sell 22,300 units of product.

Required: Determine Vernon’s margin of safety expressed as a percentage. (Round your answer to 2 decimal places. (i.e., .2345 should be entered as 23.45))
Answer #1

Contribution margin=Selling price-Variable cost

=(33-24)=$9 per unit

Hence breakeven sales=Fixed cost/Contribution margin

=(72900/9)=8100 units

Hence margin of safety=Total sales-Breakeven sales

=(22300-8100)=14200 units

Hence margin of safety%=margin of safety/Total sales

=(14200/22300)

which is equal to

=63.68%(Approx).